# Hadron Production in Ultra-relativistic Nuclear Collisions and Finite Baryon-Size Effects

### v2

@@ III Results and Discussion
[…]
@@
We find that the theoretical curves after incorporating the effect of strangeness suppression factor fit the data quite reasonably, with $\chi^{2}$/dof values for the three cases being 0.11, 1.20, and 0.85, respectively. It is noteworthy that in some previous works [13, 69, 73, 78, 12, 85, 65, 64] the factor $\gamma_{s}$ is found to increase with $\sqrt{s_{NN}}$ while in the present case we find that it decreases with $\sqrt{s_{NN}}$. The ansatz in Eq. (13) takes this aspect into account. This difference seems to emerge from the fact that in the previous analysis the hadrons were treated as purely point-like particles while in our case we have taken into account their finite sizes resulting from the baryonic hard-core repulsion. It thus indicates that the consideration of the baryonic (antibaryonic) finite sizes, leading to excluded volume type effect, play an important role in determining the values of the parameters of thermal models used for analyzing the systems produced in URHIC. The treatment of the hard-core radii of the baryons and antibaryons as a free parameter helps in understanding this aspect of the HRG.

### v3

@@ III Results and Discussion
With $\gamma_{s}^{o}$ $=$ 0.75, $y$ $=$ 0.34 GeV-1 and $z$ $=$ 11.5 GeV. The curve showing the $\sqrt{s_{NN}}$ dependence of $\gamma_{s}$ is shown in Fig. 9. We find that the factor $\gamma_{s}$ is more suppressed towards the higher collision energies. Using this approach, we redefine the theoretically obtained values of $K^{-}/\pi^{-}$, $K^{+}/\pi^{+}$, $\Xi^{+}/\pi^{-}$ as $\gamma_{s}(K^{-}/\pi^{-})$, $\gamma_{s}(K^{+}/\pi^{+})$, $\gamma_{s}^{2}(\Xi^{+}/\pi^{-})$. In this approach a factor of $\gamma_{s}$ is multiplied for each strange quark (antiquark) contained in a given hadronic specie in the system. We have demonstrated the impact of strangeness suppression factor on the $K^{-}/\pi^{-}$ ratio in Fig. 10. The green line represents a curve without $\gamma_{s}$, while the blue line depicts the results after taking into account this factor i.e., $\gamma_{s}$. It is evident that the experimental data only aligns with the model results after considering this factor. Hence we have applied this factor in all strange-to-nonstrange ratios in this calculation. The theoretical results of these ratios thus obtained after this correction emanating as a consequence of the partial saturation of the strange phase space are shown in Figs. 11 and 12.
@@
We find that the theoretical curves after incorporating the effect of strangeness suppression factor fit the data quite reasonably, with $\chi^{2}$/dof values for the three cases being 0.11, 1.20 and 0.85 respectively. It is noteworthy that in some previous works [13, 68, 72, 77, 12, 84, 64, 63] the factor $\gamma_{s}$ is found to increase with $\sqrt{s_{NN}}$ while in the present case we find that it decreases with $\sqrt{s_{NN}}$. This is because of the finite size effects that we include in this calculation while in previous works mentioned above have not incorporated this effect. The ansatz in Eq. (12) takes this aspect into account. This difference seems to emerge from the fact that in the previous analysis the hadrons were treated as purely point-like particles while in our case we have taken into account their finite sizes resulting from the baryonic hard-core repulsion. It thus indicates that the consideration of the baryonic (antibaryonic) finite sizes, leading to excluded volume type effect, play an important role in determining the values of the parameters of thermal models used for analyzing the systems produced in URHIC. The treatment of the hard-core radii of the baryons and antibaryons as a free parameter helps in understanding this aspect of the HRG.